# CYCLO: Cyclic Graph Transformer Approach to Multi-Object Relationship Modeling in Aerial Videos

### v1

@@ 5.3 Comparisons with Baseline Methods
Table 4 shows CYCLO outperforms other methods (Section B.3) across metrics and tasks, surpassing Transformer by 2.95%, 3.03%, and 2.44% in the PredCls, SGCls, and SGDet at R@20, and maintaining its lead even at higher Recall thresholds. Moreover, Table 5 shows significant improvements at mR@20, mR@50, and mR@100, with CYCLO leading HIG by 0.86% in the PredCls task at mR@20 and exceeding HIG by 1.12% and 1.32% in the SGCls and SGDet tasks, respectively, at mR@20. In addition, the consistent performance and low standard deviation of the CYCLO model across Table 4 and  5 demonstrate its robustness and overall superiority. Fig. 7 displays that CYCLO can capture the evolving relationships between objects by updating their positions and interactions.
@@ 5.4 Comparisons with State-of-the-Art Methods
Figure 7: Scene graphs generated by the CYCLO model on the AeroEye dataset. (Best viewed in colors)
@@
Performance on PVSG. Table 6 shows that the CYCLO model outperforms other models on the PVSG dataset, achieving the highest Recall and mean Recall. CYCLO surpasses HIG and Transformer by 1.23% and 1.81% at R@20, respectively, and outperforms them in R@50 and R@100. CYCLO also slightly improves mean Recall, demonstrating its robustness in periodic actions in PVSG (e.g., cooking, washing).
Table 7: Comparative performance (%) of our model and previous methods on the ASPIRe dataset.
@@
Performance on ASPIRe. The ASPIRe dataset comprises five interactivity types; however, we focus on position and relation for a fair comparison. Table 7 shows that CYCLO consistently outperforms existing methods across varying recall and mean recall thresholds. Regarding R@20, our model higher 0.69% and 5.03% than HIG for position and relation. Moreover, CYCLO significantly improves the mean Recall at every threshold, revealing that our proposed approach can address the long-tail distribution of the ASPIRe dataset.
@@ 6 Conclusions
We have presented CYCLO, an approach that captures periodic and overlapping relationships, handles long sequences, and minimizes information loss. We also introduced AeroEye, a diverse dataset of drone-captured scenes with complex object relations and positions. Extensive experiments on AeroEye, ASPIRe, and PVSG datasets demonstrate the effectiveness of our approach.
Limitations. Although CYCLO has achieved impressive performance, it may reveal limitations when dealing with incomplete or discontinuous videos. The periodic and cyclic attention mechanisms, crucial for capturing temporal and spatial object relationships, heavily rely on video continuity and completeness. Interruptions in the sequence, such as missing or discontinuous frames, disrupt the formation of accurate cyclical references, leading to inconsistent and incorrect predictions.
Broader Impacts. The proposed approach advances VidSGG which enables better capture of object interactions and temporal evolution in aerial videos, which is crucial for surveillance, disaster response, traffic management, and precision agriculture applications. However, it is important to consider the potential misuse of the proposed approach, such as unauthorized surveillance.
@@ A.1 Relationship Definition
Our AeroEye dataset focuses on capturing the spatial positions and relationships between objects, including person-to-object and object-to-object interactions. Table A.8,  A.9 provide a comprehensive list of the defined positions and relationships. Furthermore, Table A.10 offers guidance on the final relationship vocabulary specific to each scene category within the dataset.
@@
capturing exploring wildlife habitat parallel teamates
@@
celebrating farmer and harvesting tool participating in festivities towarding destination
@@
coaching gathering polices under surveillance
@@ A.2 Annotation Pipeline
    • All bounding box annotations are performed manually by skilled annotators, without relying on automated detection or tracking models. This ensures precise object localization tailored to the specific characteristics of aerial videos.
@@ A.3 Data Format
Table A.11: Comparison of available datasets for scene graph generation. The top two blocks present image and video scene graph datasets, while the next two focus on image and drone video datasets. Viewpoints include five types: from ego (1st-person) view to 3rd-person view, and drone-captured perspectives, which is our main focus in this work (/ in colors), including aerial (top-down), oblique (slanted), and ground (eye-level) perspectives. # denotes the number of corresponding items. The best values in drone blocks are highlighted.
BBox Relation #Instances 3rd ego aerial oblique ground
Visual Genome [34] - 108K - 3.8M 33K 42K -
VG-150 [64] - 88K - 2.8M 150 50 -
VrR-VG [65] - 59K - 282.4K 1.6K 117 -
GQA [66] - 85K - 1.7K 310 -
PSG [37] - 49K - 538.2K 80 56 -
VidVRD[67] 1K 296.2K 1920 $\times$ 1080 15.1K 35 132 -
Action Genome [3] 10K 234.3K 1280 $\times$ 720 476.3K 25 25 -
ASPIRe [5] 1.5K 1.6M 1280 $\times$ 720 167.8K 833 4.5K 7
SportsHHI [35] 80 11.4K 1280 $\times$ 720 118.1K 1 34 2
VidOR [68] 10K 55.4 640 $\times$ 360 50K 80 50 1
VidSTG [69] 10K 55.4 640 $\times$ 360 50K 80 50 1
EPIC-KITCHENS [70] 700 11.5K 1920 $\times$ 1080 454.3K 21 13 1
PVSG [4] 400 153K 1920 $\times$ 1080 7.6K 126 57 7
DOTA [13] - 11.3K 1490 $\times$ 957 1.8M 18 - -
AI-TOD [20] - 28.1K 800 $\times$ 800 700.6K 8 - -
DIOR-R [21] - 23.5K 800 $\times$ 800 192.5K 20 - -
MONET [22] - 53K - 162K 3 - -
SODA-A [23] - 2.5K 4761 $\times$ 2777 872.1K 9 - -
VisDrone [14] 288 261.9K 3840 $\times$ 2160 2.6M 10 - -
UAVDT [24] 100 40.7K 1080 $\times$ 540 0.84M 3 - 6
Stanford Drone [25] 10K 929.5K - - 6 - 8
UIT-ADrone [26] 51 206.2K 1920 $\times$ 1080 210.5K 8 - 1
ERA [27] 2.9K 343.7K 640 $\times$ 640 - - - 25
MOR-UAV [28] 30 10.9K 1920 $\times$ 1080 89.8K 2 - -
AU-AIR [29] - 32.8K 1920 $\times$ 1080 132K 8 - -
DroneSURF [30] 200 411.5K 1280 $\times$ 720 786K 1 - -
MiniDrone [31] 38 23.3K 224 $\times$ 224 - - - 1
Brutal Running [32] - 1K 227 $\times$ 227 - - - 1
UAVid [33] 30 300 4096 $\times$ 2160 - 8 - -
MAVREC [15] 24 537K 3840 $\times$ 2160 1.1M 10 - 4
AeroEye (Ours) 2.3K 261.5K 3840 $\times$ 2160 1.8M 57 384 29
@@ A.5 Data Samples
Appendix B Methodology
B.1 Transformer
The Transformer [57] is a neural network architecture consisting of stacked layers, each employing multi-head attention. Given an input matrix $X\in\mathbb{R}^{N\times D}$, with $N$ entries and dimensionality $D$, the transformer linearly projects $X$ into queries ($Q$), keys ($K$), and values ($V$) using learned weights $W_{Q},W_{K},W_{V}$, respectively. Self-attention is computed by:
$\text{Attention}(Q,K,V)=\text{Softmax}\left(\frac{QK^{T}}{\sqrt{D_{k}}}\right)V,$ (B.6)
where $D_{k}$ is the dimensionality of the keys. Multi-head attention extends this by splitting $Q$, $K$, and $V$ into multiple subspaces, enabling the model to simultaneously attend to different representation subspaces. Each head $h_{i}$ is computed as:
$h_{i}=\text{Attention}(XW_{Q_{i}},XW_{K_{i}},XW_{V_{i}}).$ (B.7)
The head outputs are concatenated and linearly transformed:
$\text{MultiHead}(Q,K,V)=\text{Concat}(h_{1},\dots,h_{e})W_{O}.$ (B.8)
Finally, the multi-head attention output is normalized, passed through a feed-forward network, and normalized again with a residual connection to form a complete transformer layer.
B.2 Concepts
B.2.1 Progression
@@
B.2.2 Batch Progression
In contrast to the Progressive approach, the Batch Progressive approach passes the relationship features through a Transformer architecture (detailed in Appendix B.1) before classification.
@@ B.2.3 Hierarchical Graph
B.3 Baseline Methods
In the experiments shown in Table 6 and Table 7, we compare our method with four baseline approaches introduced in [37]. The Vanilla approach, also known as the Progressive approach, and the 1D Convolution, Handcrafted, and Transformer approaches are all considered Batch Progressive methods. The Vanilla and Transformer approaches are described in more detail in Section B.2. The Handcrafted and 1D-Convolutional approaches are introduced below:
Handcrafted. To incorporate temporal information effectively, the handcrafted filter is designed to integrate context from adjacent frames, enhancing the temporal coherence in feature representation. The filter (e.g., Gaussian), applied across the temporal dimension of the feature vectors:
$h_{t}^{k}=\sum_{i=-W/2}^{W/2}g_{i}\cdot x_{t+i}^{k}$ (B.13)
where $x_{t}^{k}$ represents the feature vector, $W$ is the kernel size, and $w_{i}$ are the learnable weights. The weights $w_{i}$ are defined by the vector $[1/4,1/2,1,1/2,1/4]$ with a window size of 5 frames.
1D-Convolutional. A learnable 1D-Convolutional layer uses CNN consisting of three 1D convolutional layers with a kernel size of 5, allowing the network to learn optimal weights for combining features across frames. The convolutional filter are defined as:
$h_{t}^{k}=\sum_{i=-W/2}^{W/2}w_{i}\cdot x_{t+i}^{k}$ (B.14)
where $x_{t}$ is the input feature, $W$ is the kernel size, and $w_{k}$ are the learnable weights.
B.4 Properties of the CYCLO method
B.4.1 Spatial Attention Graph
Figure B.12: Illustration of the Attention Spatial Graph, where a static graph is extracted from each frame of an input video. The nodes in the graph represent spatial regions, and the edges represent the attention-based relationships between these regions.
The CYCLO model computes spatial relationships through Eqn. (1) and (2) at each layer $l$ and for each frame $t$, as illustrated in Fig.B.12. This computation analyzes interactions between query and key vectors through an attention mechanism, allowing CYCLO to explicitly capture and dynamically update the relationships between objects across different layers.
The primary distinction between CYCLO and the previous approach (refer to Section 2.2 and Section B.2) lies in their mechanisms for establishing and refining object relationships in each frame. CYCLO computes relational representations for each object pair at every layer, providing a detailed and adaptable understanding of object interactions. CYCLO also enhances SGG through multi-head self-attention layers in transformer-based architectures, where each “head” focuses on different facets of relationships, yielding a comprehensive understanding of scene dynamics. In addition, the direct computation in CYCLO eliminates the need for a separate relational detection phase, reducing computational overhead and error propagation from sequential processing. Conversely, the Batch Progressive approach employs a unified self-attention mechanism within the Transformer Encoder (Eqn. (B.10)) to iteratively merge and refine relationship features.
B.4.2 Circle and Periodic
Fourier Transform. The Fourier transform is a fundamental that enables the analysis and representation of periodic functions as a sum of sinusoidal components. Given a periodic function $f(t)$ with period $T$, the Fourier series representation can be expressed as:
$f(t)=\sum_{n=-\infty}^{\infty}c_{n}e^{i2\pi\frac{nt}{T}},$ (B.15)
where the complex coefficients $c_{n}$ are computed via the integral:
[…]
The Fourier series representation in Eqn.(B.15) expresses a periodic function as an infinite sum of complex exponentials, each representing a specific frequency component. The coefficients $c_{n}$ in this representation, computed using the integral in Eqn.(B.16), capture the amplitude and phase of each frequency component. This decomposition of a periodic function into its constituent frequencies explains its behavior and properties in the frequency domain.
Cyclic Attention. We leverage the Fourier transform for sequence analysis in cyclic attention (CA), which considers a periodic sequence $K$ with period $T$. Inspired by the Fourier series representation, we can express the sequence $K$ as:
$K(t)=\sum_{n=-\infty}^{\infty}c_{n}e^{i2\pi\frac{nt}{T}},$ (B.17)
where the coefficients $c_{n}$ capture the frequency-domain characteristics of the sequence. The modulo operation in CA ensures proper handling of cyclic shifts, as shown in the following equation:
$(t+i)\ \mathrm{mod}\ T=\begin{cases}t+i-T&\text{if }t+i\geq T,\\ t+i&\text{if }t+i<T.\end{cases}$ (B.18)
By applying the modulo operation to the Fourier series representation of $K$, we obtain:
$K_{(t+i),\mathrm{mod},T}=\sum_{n=-\infty}^{\infty}c_{n}e^{i2\pi\frac{n(t+i)}{T}}.$ (B.19)
Then, we incorporate Eqn. (B.19) into the CA mechanism yields:
$\text{CA}(Q_{t},K_{t})=\sum_{i=0}^{T-1}\sigma\left(\frac{Q_{t}\left(\sum_{n=-\infty}^{\infty}c_{n}e^{i2\pi\frac{n(t+i)}{T}}\right)^{\top}}{\sqrt{d_{\text{head}}}}\right)$ (B.20)
Eqn. (B.20) reveals the profound connection between cyclic attention and the Fourier transform. Our cyclic attention evaluates the interaction between the query $Q_{t}$ and cyclically shifted versions of the key $K_{t}$, with each shift weighted by the corresponding Fourier series exponential term. The summation over $i$ captures the cumulative influence of all cyclic shifts, providing a comprehensive assessment of the sequence’s periodicity. In addition, by leveraging the inherent periodicity of sequences, cyclic attention effectively exploits the underlying structure analogous to how the Fourier series decomposes a function into its constituent frequencies. Each term in the summation of Eqn. (B.20) represents a specific phase shift in the Fourier series, establishing a direct link between the time-domain interactions in cyclic attention and the frequency-domain analysis in Fourier theory. Furthermore, this mathematical perspective illuminates the ability to capture and utilize the rich information embedded in the sequence’s periodic structure of cyclic attention.
B.4.3 Non-Permutation Equivariance
Self-Attention and Permutation Equivariance. Different from Eqn. (3), the computation for self-attention can be defined as:
$\text{SA}(Q_{t},K_{t})=\text{softmax}\left(\frac{Q_{t}K_{t}^{\top}}{\sqrt{d_{\text{head}}}}\right)V_{t}$ (B.21)
When we apply a permutation matrix $\pi$ to $Q_{t}$ and $K_{t}$, we obtain the transformed matrices ${Q^{\prime}}_{t}=Q_{t}\pi$ and ${K^{\prime}}_{t}=K_{t}\pi$. Given that the permutation matrix is orthogonal, $\pi^{\top}=\pi^{-1}$, we have:
$\text{SA}({Q^{\prime}}_{t},{K^{\prime}}_{t})$ $=\text{softmax}\left(\frac{{Q^{\prime}}_{t}{K^{\prime}}_{t}^{\top}}{\sqrt{d_{\text{head}}}}\right)V_{t}$ (B.22)
$=\text{softmax}\left(\frac{Q_{t}\pi(\pi^{\top}K_{t}^{\top})}{\sqrt{d_{\text{head}}}}\right)V_{t}$ (B.23)
$=\text{softmax}\left(\frac{Q_{t}K_{t}^{\top}}{\sqrt{d_{\text{head}}}}\right)V_{t}=\pi(\text{SA}(Q_{t},K_{t}))$ (B.24)
Eqn. (B.22) demonstrates that self-attention is permutation equivariant, as the output under permutation is simply a permuted version of the original output.
Cyclic Attention and Non-Permutation Interaction. We consider the permutation $\pi$ applied to $Q_{t}$ and $K_{t}$ in Eqn. (3), leading to ${Q^{\prime}}_{t}=Q_{t}\pi$ and ${K^{\prime}}_{t}=K_{t}\pi$:
$\text{CA}({Q^{\prime}}_{t},{K^{\prime}}_{t})$ $=\text{softmax}\left(\frac{{Q^{\prime}}_{t}[({K^{\prime}}_{t})_{(t+i)\mod T}]^{\top}}{\sqrt{d_{\text{head}}}}\right)V_{t}$ (B.25)
$=\text{softmax}\left(\frac{Q_{t}\pi[(K_{t}\pi)_{(t+i)\mod T}]^{\top}}{\sqrt{d_{\text{head}}}}\right)V_{t}$ (B.26)
In Eqn. (B.25), the key difficulty arises in expressing $(K_{t}\pi)_{(t+i)\mod T}$ due to the cyclic shift applied after the permutation. The indices affected by the cyclic shift might no longer align with those modified by the permutation, leading to:
$(K_{t}\pi)_{(t+i)\mod T}\neq(K_{(t+i)\mod T}\pi)$ (B.27)
The discrepancy in Eqn. (B.27) underscores the difference between self-attention and cyclic attention regarding permutation equivariance. While self-attention treats elements based solely on their content, resulting in a permutation equivariant output, cyclic attention indicates a complex interaction between cyclic shifts and permutations. This interaction renders the output of cyclic attention under non-permutation, as the relative order of elements is altered in a way that is not simply a permutation of the original output. Consequently, cyclic attention is sensitive to the intrinsic order and continuity of the sequence, enabling it to capture the temporal relationships between objects. This property makes cyclic attention particularly well-suited for tasks where the sequential nature of the data is essential, as it can effectively model the underlying structure and dependencies present in the sequence.

### v4

@@ 5.3 Comparisons with Baseline Methods
Table 4 shows CYCLO outperforms other methods across metrics and tasks, surpassing Transformer by 2.95%, 3.03%, and 2.44% in the PredCls, SGCls, and SGDet at R@20, and maintaining its lead even at higher Recall thresholds. Moreover, Table 5 shows significant improvements at mR@20, mR@50, and mR@100, with CYCLO leading HIG by 0.86% in the PredCls task at mR@20 and exceeding HIG by 1.12% and 1.32% in the SGCls and SGDet tasks, respectively, at mR@20. In addition, the consistent performance and low standard deviation of the CYCLO model across Table 4 and  5 demonstrate its robustness and overall superiority. Fig. 7 displays that CYCLO can capture the evolving relationships between objects by updating their positions and interactions.
@@ 5.4 Comparisons with State-of-the-Art Methods
We compare our performance on two recent benchmark datasets on VidSGG (i.e. PVSG and ASPIRe).
@@
Table 7: Comparative performance (%) of our model and previous methods on the ASPIRe dataset, evaluated by Recall (R) and mean Recall (mR).
@@
Performance on PVSG. In Table 7, we report that the CYCLO approach achieves superior performance compared to other models on the PVSG dataset, particularly in terms of Recall and mean Recall metrics. In particular, CYCLO surpasses HIG and Transformer by 1.23% and 1.81% at R@20, respectively. In addition, it consistently outperforms both models at higher recall rates, including R@50 and R@100, showcasing its effectiveness across various thresholds. Our approach also slightly improves the mean Recall, as the experimental results demonstrate. These results highlight the robustness of our CYCLO approach in recognizing and handling periodic actions, such as cooking, washing, cleaning, exercising, and other routine tasks frequently represented on the PVSG dataset.
Performance on ASPIRe. The ASPIRe dataset includes five distinct interactivity types. However, we focus on position and relation to ensure a fair comparison. As shown in Table 7, CYCLO consistently outperforms existing models across multiple recall and mean recall thresholds. Notably, at R@20, our CYCLO approach outperforms the HIG method, the second-best model, by 0.69% in position and a more substantial 5.03% in relation. Additionally, CYCLO shows significant gains in mean Recall across all evaluated thresholds, demonstrating its effectiveness in tackling the long-tail distribution.
Figure 7: Scene graphs generated by the CYCLO model on the AeroEye dataset, illustrating dynamic relationships between objects and agents across UAV-captured frames. (Best viewed in colors)
@@ 6 Conclusions
We have introduced CYCLO, a novel approach that effectively captures periodic and overlapping relationships, handles extended sequences, and minimizes information loss, making it suitable for complex temporal modeling. In addition, we presented AeroEye, a comprehensive and diverse dataset composed of drone-captured scenes, specifically designed to represent intricate object relationships and spatial positions in aerial videos. Through extensive experiments on the AeroEye dataset and two large-scale in-the-wild datasets (i.e. ASPIRe and PVSG), we demonstrated the robustness and effectiveness of CYCLO in capturing dynamic interactions and evolving relationships over time.
Limitations. Although our CYCLO approach has achieved impressive performance, it may reveal limitations when dealing with incomplete or discontinuous videos. The periodic and cyclic attention mechanisms, crucial for capturing temporal and spatial object relationships, heavily rely on video continuity and completeness. Interruptions in the sequence, such as missing or discontinuous frames, disrupt the formation of accurate cyclical references, leading to inconsistent and incorrect predictions.
Broader Impacts. The proposed approach improves the capture of object interactions and temporal evolution in aerial and in-the-wild videos, which is critical for surveillance, disaster response, traffic management, and precision agriculture applications. By modeling object interactions over time, CYCLO supports more informed decision-making, leading to safer and more sustainable practices. This advancement opens the door for future work developing surveillance systems that can model complex relationships from drone videos. However, it is important to recognize the potential risks associated with this approach, particularly the possibility of using it for unauthorized surveillance.
Acknowledgment. This work is partly supported by J.B. Hunt Transport Services (JBHunt), NSF Data Science and Data Analytics that are Robust and Trusted (DART), NSF SBIR Phase 2, and Arkansas Biosciences Institute (ABI) grants. We also acknowledge Thanh-Dat Truong for invaluable discussions and the Arkansas High-Performance Computing Center (AHPCC) for providing GPUs.
@@ A.1 Relationship Definition
Our AeroEye dataset focuses on capturing the spatial positions and relationships between objects, including person-to-object and object-to-object interactions. Table A.8,  A.9 provides a comprehensive list of the defined positions and relationships. Furthermore, Table A.10 offers guidance on the final relationship vocabulary specific to each scene category within the dataset.
@@
capturing exploring wildlife habitat parallel teammates
@@
celebrating farmer and harvesting tool participating in festivities toward destination
@@
coaching gathering police under surveillance
@@ A.2 Annotation Pipeline
    • All bounding box annotations are performed manually by skilled annotators without relying on automated detection or tracking models. This ensures precise object localization tailored to the specific characteristics of aerial videos.
@@ A.5 Data Samples
Appendix B Concepts
B.1 Progression
@@
B.2 Batch Progression
In contrast to the Progressive approach, the Batch Progressive approach passes the relationship features through a Transformer architecture before classification.